# Deep learning-derived arterial input function

### v1

@@ 1 Introduction
Dynamic positron emission tomography (PET) is a molecular imaging technique involving the acquisition of sequential PET images over time, following radiotracer injection. By analyzing these dynamic PET data through kinetic modeling, one can quantify key physiological and biochemical parameters, including tissue receptor density, tracer influx or trapping rates, and competitive interactions between endogenous and exogenous ligands [13, 48]. Consequently, dynamic PET coupled with kinetic modeling has emerged as a crucial approach for investigating various diseases, particularly neurological disorders.
@@
In kinetic modeling, sets of ordinary differential equations (ODEs) describe the dynamic relationships between a radiotracer, its physiological states, and the resulting PET images [21]. These tracer-specific models typically originate from compartmental frameworks in which the plasma concentration of the radiotracer over time serves as the input function [64]. The coefficients within these ODEs are solved using PET data, capturing the intrinsic kinetic properties of the tracer. Accurate determination of this input function is critical for robust kinetic modeling. Traditionally, this involves serial arterial blood sampling from the patient’s radial artery, beginning at tracer injection (the start of dynamic PET scanning) and continuing until the scan concludes (as shown in the left panel of Fig. 1). Collected blood samples are centrifuged to isolate plasma, and the radiotracer concentration is measured using a well counter, with corrections applied for radiolabeled metabolites, thus producing a metabolite-corrected AIF [8]. Although serial arterial sampling is considered reasonably safe, dynamic PET scans typically last 90 to 180 minutes, and patient tolerance to arterial sampling can vary significantly, especially among elderly or medically compromised populations. This variability may lead to slower recruitment or increased dropout rates in longitudinal studies. Moreover, risks associated with arterial cannulation rise in aging patients, particularly those on medications that affect blood coagulation, potentially contraindicating arterial line placement [25]. Consequently, such concerns may deter researchers from adopting promising radiotracers in multicenter, longitudinal trials due to reluctance or the inability to perform arterial sampling. In addition, some PET research facilities may not have the capability to readily acquire dynamic, arterial measurements.
@@
Previous research has explored alternative, non-invasive methods for deriving the AIF, primarily focusing on population-based input functions (PBIF) and image-derived input functions (IDIF) [66]. PBIF generates a standardized input function by averaging and normalizing arterial data across multiple subjects. However, this approach does not account for individual physiological differences, variations in scanner characteristics, or the quality of image acquisition and reconstruction algorithms, leading to a notable limitation due to its generalized nature [66, 2]. Additionally, PBIF cannot effectively distinguish between the parent radiotracer and its radiolabeled metabolites in the blood, typically requiring supplementary blood samples to perform accurate metabolite corrections [65, 56]. In contrast, IDIF provides a more personalized method by directly estimating the whole-blood time-activity curve from an individual’s dynamic PET images. By accommodating patient-specific variations, IDIF can yield more accurate and individualized estimates. IDIF techniques are generally categorized into segmentation-based and statistical decomposition-based approaches. Despite progress in IDIF estimation methods, significant challenges have impeded their broader clinical adoption [65]. Major obstacles include the need for precise carotid artery segmentation, reliance on blood sampling at specific time points for calibration, and the difficulty of distinguishing the parent radiotracer from its radiolabeled metabolites.
@@
Recently, deep learning has emerged as a promising technique across various areas of medical imaging. Although still in its early stages, initial studies integrating deep learning methods have demonstrated improved accuracy AIF estimation, surpassing traditional IDIF and PBIF approaches [61, 18, 12, 6]. Our group previously pioneered the use of deep neural networks (DNNs) to directly estimate AIF from PET images [61]. In this approach, a DNN processes two 3D PET volumes—one at a particular time point and another averaged over all time points—to predict the AIF value at each moment and subsequently reconstruct the full AIF curve. However, the method has limitations: since the network does not incorporate the entire dynamic PET sequence, it occasionally produces suboptimal estimates with non-smooth tails. To address this limitation, Ferrante et al. recently introduced a physically informed neural network (PINN), which estimates parameters of an analytical AIF model represented by a combination of two Gaussians and an exponential term modulated by a sigmoid function [18]. Alternatively, Cui et al. bypassed AIF estimation entirely, directly reconstructing parametric images using an unsupervised deep learning framework known as the conditional deep image prior [12]. Nevertheless, these innovative methods remain constrained by limited sample sizes—stemming from inherent blood sampling difficulties—and by relatively low robustness when applied across different radiotracers.
@@
In this study, we aim to overcome the limitations of current AIF estimation techniques and eliminate the need for manual intervention by developing a deep-learning-based method termed the deep-learning-derived input function (DLIF). Building upon our preliminary work [61, 6], DLIF directly estimates metabolite-corrected AIF from dynamic brain PET data. Once trained, DLIF provides a fully non-invasive alternative, circumventing the challenges associated with traditional image segmentation methods and invasive arterial blood sampling. This advancement promises immediate benefits to ongoing and future research by improving patient comfort, significantly cutting operational costs related to anesthesiologist services, blood sampling, personnel, and facility logistics. Additionally, DLIF has the potential to enhance participant recruitment and retention in longitudinal studies.
@@ 2.1 Segmentation-based IDIF methods
Segmentation-based IDIF relies on identifying and segmenting large blood vessels, such as segments of the aorta or femoral arteries, to estimate the whole-blood time-activity curve for individual patients. These vessels are chosen primarily due to their large size, which facilitates effective correction of partial volume effects (PVE) [66]. However, in brain PET studies, the small size of cerebral vessels often results in significant PVE, substantially decreasing the accuracy and precision of the derived IDIFs [66]. To address this challenge, numerous segmentation methods have been proposed, utilizing either PET images alone or PET images co-registered with MRI [31, 7, 60, 30, 44, 54, 38, 20, 28, 19, 53, 26]. To further mitigate the combined impacts of PVE and noise, statistical decomposition-based approaches have been adopted for IDIF estimation from segmented data. For example, Chen et al. proposed a method that involves manually delineating the carotid arteries on dynamic PET images [7]. Their approach explicitly accounts for PVE by modeling the signal from each voxel within the carotid artery region-of-interest (ROI) as a mixture of true vascular radioactivity and spillover radioactivity from adjacent tissues. The relationship is mathematically expressed as:
@@
Building upon this approach, Su et al. replaced invasive arterial blood sampling with an independent component analysis (ICA)-based method applied to the initial 30 minutes of dynamic PET data [54]. They also introduced automation by applying Gaussian thresholding to segment arteries and dilation methods to isolate adjacent tissues. Parker et al. further enhanced IDIF estimation accuracy by incorporating principal component analysis (PCA) for dimensionality reduction and employing a graph-based Mumford-Shah segmentation algorithm for improved delineation of carotid arteries [44]. More recent methods have continued to advance artery segmentation by integrating anatomical imaging modalities and adopting sophisticated segmentation techniques [20, 26, 55, 53, 19].
@@ 2.2 Direct estimation of IDIF
Alternatively, IDIF can also be estimated directly from voxel values in dynamic PET, avoiding explicit segmentation of the carotid arteries [40, 41, 42, 3, 63]. Such approaches commonly rely on statistical decomposition techniques, particularly independent component analysis (ICA), to separate voxel signals into plasma and tissue components. However, these methods typically cannot fully eliminate invasive blood sampling; usually, at least one blood sample is required at a specific time point to calibrate the estimated IDIF. For example, Naganawa et al. proposed the following model to describe voxel values in dynamic PET images [40, 41]:
@@
where $s_{p}(q)$ and $s_{t}(q)$ represent the plasma and tissue contributions of the voxel $q$, respectively, while $c_{p}(t)$ and $c_{t}(t)$ denote the uncalibrated IDIF and the tissue time activity curve (tTAC) at time $t$, respectively. When expressed in matrix form, Eqn. 2 transforms into $\mathbf{X}=\mathbf{CS}$, where a modified ICA method known as EPICA decomposes the dynamic PET data matrix $\mathbf{X}$ into component matrices $\mathbf{C}$ and $\mathbf{S}$. The plasma-related component in $\mathbf{S}$ provides the uncalibrated IDIF. Since ICA cannot inherently determine the scale and sign of its estimated components, a single blood sample, typically at the peak of the true AIF, is obtained for calibration purposes. Similarly, Bodvarsson et al. employed non-negative matrix factorization (NMF) instead of ICA to perform voxel decomposition and derive the IDIF [3]. Building upon previous approaches, Naganawa et al. later enhanced the intersectional searching algorithm (ISA) [63], developing EPISA. EPISA estimates a time-integrated IDIF directly from PET images suitable for Logan graphical analysis [35], utilizing distinct tTACs derived from image regions grouped by intensity-based clustering methods.
@@ 3 Methods
$\begin{split}I(t,\boldsymbol{q})&=f_{b}(\boldsymbol{q})\Big{[}\alpha(\boldsymbol{q})c_{p}(t)+\beta(\boldsymbol{q})c_{r}(t)\Big{]}+f_{t}(\boldsymbol{q})c_{t}(t)\\ \end{split}$ (3)
@@
To achieve our goal, we introduce a novel DNN framework. This network is built on top of ViT [15], enhanced with additional modules customized to process time-sequenced image volumes, such as dynamic PET scans. The output of ViT consists of a set of parameters that define the shapes of basis functions, which are then used to model the DLIFs.
@@ 3.1.1 Vision Transformer
Transformers, initially developed for natural language processing tasks [58], have demonstrated significant potential in computer vision tasks [15, 33, 22]. Following these successes, they have been increasingly applied in the processing of medical images [29]. A notable advantage of Transformers is their scalability [67, 34] and their capacity to capture long-range relationships between parts of the input. The decision to choose ViT over convolutional neural networks (ConvNets) for our application of AIF estimation is based on clear reasoning. ConvNets typically adopt small convolution kernels, usually ranging in size from 3 to 7, which inherently assume strong correlations between adjacent points (i.e., inductive bias). This presents two drawbacks for our application. First, AIF estimation often involves comparing spatially distant voxels, as seen in ISA [63] and EPICA [42] described in the Introduction section. Localized convolution kernels in ConvNets are not as effective in capturing these distant relationships, which only become implicit as the network layers deepen. Second, for precise estimation across the temporal dimension, it is preferable to process the entire temporal data directly, rather than applying small convolution kernels across time points. It is important to acknowledge recent advancements that show that ConvNets can be optimized with much larger kernels [14, 32] or advanced convolutional operations [34], potentially matching the capabilities of Transformers. However, optimizing the network architecture to balance Transformers and ConvNets is beyond the scope of our study. In this work, we have benchmarked the proposed network against commonly used DNNs, with further details and results presented in the Results section.
The proposed network is built on the foundation of ViT [15], which was originally designed for 2D image processing. Recognizing the limitations of ViT in this regard, we have expanded its functionality to effectively handle 3D image volumes. As depicted in Fig. 2, the downsampled dynamic PET image $I\in\mathbb{R}^{T\times\frac{H}{4}\times\frac{W}{4}\times\frac{L}{4}}$ is first divided into a series of $N$ non-overlapping patches $\{I_{1},I_{2},...,I_{N}\}$, each of size $T\times P\times P\times P$, that is, $I_{i}\in\mathbb{R}^{T\times P\times P\times P}$. Subsequently, each patch is vectorized and linearly projected to create tokens with a $D$-dimensional embedding. This is formulated as:
so that $\hat{\mathbf{z}}$ assumes the shape of $N\times D$. Next, we append separate learnable embeddings/tokens ($I^{\text{spatial}}\in\mathbb{R}^{1\times D}$) to both the sequence of embedded patches and the temporal dimension, respectively. Consequently, $\hat{\mathbf{z}}_{0}$ is expanded to a size of $(N+1)\times D$. This token interacts with all other tokens across spatial dimensions, which resembles the traditional methods that compare voxels across varying spatial distances [63, 42]. In a later stage, a specifically designed estimation head will be attached to $I^{\text{spatial}}$ to generate the DLIF. Finally, before progressing to the Transformer encoder, a learnable positional embedding is added to the tokens, enabling them to preserve spatial information throughout the network. The process is described as:
$\begin{split}\hat{\mathbf{z}}_{\ell}=\text{MSA}(\text{LN}(\mathbf{z}_{\ell-1}))+\mathbf{z}_{\ell-1},\ \ \ \ &\ell=1...L,\\ \mathbf{z}_{\ell}=\text{MLP}(\text{LN}(\hat{\mathbf{z}}_{\ell}))+\hat{\mathbf{z}}_{\ell},\ \ \ \ &\ell=1...L,\\ \end{split}$ (6)
where $\text{LN}(\cdot)$ indicates layer normalization [1]. The operations $\text{MSA}(\cdot)$ and $\text{MLP}(\cdot)$ represent the multi-head self-attention mechanism and the multi-layer perceptron, respectively.
3.1.2 Self-attention
Here, we adopt the standard softmax-based self-attention operation [58], where the tokens $\mathbf{z}$ are first projected into three separate embeddings: a query $\mathbf{Q}$, a key $\mathbf{K}$, and a value $\mathbf{V}$ using a set of learnable parameters $\mathbf{U}_{qkv}$:
$\begin{split}[\mathbf{Q},\mathbf{K},\mathbf{V}]=\mathbf{z}\mathbf{U}_{qkv}&,\ \ \ \ \mathbf{U}_{qkv}\in\mathbb{R}^{D\times 3D_{h}},\\ \mathbf{A}=\text{softmax}(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{D_{h}}})&,\ \ \ \ \mathbf{A}\in\mathbb{R}^{(N+1)\times(N+1)},\\ \text{SA}(\mathbf{z})=\mathbf{A}\mathbf{V}.&\end{split}$ (7)
@@ 3.1.3 Estimation head
In the original ViT [15], the classification head is attached to the class token for the final prediction. In a similar fashion, our model attaches different estimation heads to the estimation token $I^{\text{spatial}}\in\mathbb{R}^{1\times D}$. Since this token engages with all others during self-attention, it is capable of learning the global information necessary for AIF estimation. The estimation heads are responsible for producing either a direct AIF estimation, which has an output dimension of $1\times T$, or a set of parameters to combine the basis functions. The latter is discussed in the following section.

### v2

@@ 1 Introduction
Dynamic positron emission tomography (PET) is a molecular imaging technique involving the acquisition of sequential PET images over time, following radiotracer injection. By analyzing these dynamic PET data through kinetic modeling, one can quantify key physiological and biochemical parameters, including tissue receptor density, tracer influx or trapping rates, and competitive interactions between endogenous and exogenous ligands [dimitrakopoulou2021kinetic, rahmim2019dynamic]. Consequently, dynamic PET coupled with kinetic modeling has emerged as a crucial approach for investigating various diseases, particularly neurological disorders.
@@
In kinetic modeling, sets of ordinary differential equations (ODEs) describe the dynamic relationships between a radiotracer, its physiological states, and the resulting PET images [gunn2001positron]. These tracer-specific models typically originate from compartmental frameworks in which the plasma concentration of the radiotracer over time serves as the input function [watabe2006pet]. The coefficients within these ODEs are solved using PET data, capturing the intrinsic kinetic properties of the tracer. Accurate determination of this input function is critical for robust kinetic modeling. Traditionally, this involves serial arterial blood sampling from the patient’s radial artery, beginning at tracer injection (the start of dynamic PET scanning) and continuing until the scan concludes (as shown in the left panel of Fig. 1). Collected blood samples are centrifuged to isolate plasma, and the radiotracer concentration is measured using a well counter, with corrections applied for radiolabeled metabolites, thus producing a metabolite-corrected AIF [coughlin2018distribution]. Although serial arterial sampling is considered reasonably safe, dynamic PET scans typically last 90 to 180 minutes, and patient tolerance to arterial sampling can vary significantly, especially among elderly or medically compromised populations. This variability may lead to slower recruitment or increased dropout rates in longitudinal studies. Moreover, risks associated with arterial cannulation rise in aging patients, particularly those on medications that affect blood coagulation, potentially contraindicating arterial line placement [kang2018noninvasive]. Consequently, such concerns may deter researchers from adopting promising radiotracers in multicenter, longitudinal trials due to reluctance or the inability to perform arterial sampling. In addition, some PET research facilities may not have the capability to readily acquire dynamic, arterial measurements.
@@
Previous research has explored alternative, non-invasive methods for deriving the AIF, primarily focusing on population-based input functions (PBIF) and image-derived input functions (IDIF) [Zanotti_Fregonara_2012]. PBIF generates a standardized input function by averaging and normalizing arterial data across subjects, but it overlooks individual physiology, scanner differences, and acquisition quality, limiting accuracy [Zanotti_Fregonara_2012, Boutin_2007]. Additionally, PBIF cannot effectively distinguish between the parent radiotracer and its radiolabeled metabolites in the blood, typically requiring supplementary blood samples to perform accurate metabolite corrections [zanotti2011image, Takikawa_1993]. In contrast, IDIF estimates the whole-blood time-activity curve directly from dynamic PET images, capturing patient-specific variability for more individualized results. IDIF methods generally fall into segmentation- or statistical decomposition-based categories. Nonetheless, challenges such as accurate carotid segmentation, calibration via blood samples, and metabolite correction continue to hinder clinical adoption [zanotti2011image].
@@
Recently, deep learning has emerged as a promising technique across various areas of medical imaging. Although still in its early stages, initial studies integrating deep learning methods have demonstrated improved accuracy AIF estimation, surpassing traditional IDIF and PBIF approaches [wang2020direct, ferrante2022physically, ferrante2024physically, Cui_2022, chen2023estimating, kuttner2024deep]. Our group previously pioneered the use of deep neural networks (DNNs) to directly estimate AIF from PET images [wang2020direct]. In this approach, a DNN processes two 3D PET volumes—one at a particular time point and another averaged over all time points—to predict the AIF value at each moment and subsequently reconstruct the full AIF curve. However, the method has limitations: since the network does not incorporate the entire dynamic PET sequence, it occasionally produces suboptimal estimates with non-smooth tails. In contrast, kuttner2024deep employed a DNN that uses the entire dynamic PET sequence as input to estimate AIF values at the corresponding time points. ferrante2022physically, ferrante2024physically further advanced this direction by introducing a physically informed neural network (PINN), which estimates parameters of an analytical AIF model represented by a combination of two Gaussians and an exponential term modulated by a sigmoid function. Alternatively, Cui_2022 bypassed AIF estimation entirely, directly reconstructing parametric images using an unsupervised deep learning framework known as the conditional deep image prior. Nevertheless, these innovative methods remain constrained by limited sample sizes—stemming from inherent blood sampling difficulties—and by relatively low robustness when applied across different radiotracers.
@@
In this study, we aim to overcome the limitations of current AIF estimation techniques and eliminate the need for manual intervention by developing a deep-learning-based method termed the deep-learning-derived input function (DLIF). Building upon our preliminary work [wang2020direct, chen2023estimating], DLIF directly estimates metabolite-corrected AIF from dynamic brain PET data. Although conceptually related to the recent work by Ferrante et al. [ferrante2022physically, ferrante2024physically], which estimates analytical parameters of the Parker model, the proposed DLIF framework offers greater modeling flexibility by representing the input function through a composition of basis functions rather than enforcing a parametric functional form. This design enables DLIF to capture complex temporal dynamics beyond those expressible by a limited number of Gaussian or exponential terms. Once trained, DLIF provides a fully non-invasive alternative, circumventing the challenges associated with traditional image segmentation methods and invasive arterial blood sampling. This advancement promises immediate benefits to ongoing and future research by improving patient comfort, significantly cutting operational costs related to anesthesiologist services, blood sampling, personnel, and facility logistics. Additionally, DLIF has the potential to enhance participant recruitment and retention in longitudinal studies.
The remainder of the paper is organized as follows. Section 2 discusses related work. Section 3 describes the proposed methodology. The experimental setup, implementation details, and datasets used in this study are discussed in Sect. 4. Section 5 presents the experimental results. The findings drawn from these results are discussed in Sect. 6, and Sect. 7 concludes the paper. For clarity, all abbreviations used in this paper are summarized in Appendix 5.
@@ 2.1 Segmentation-based IDIF methods
Segmentation-based IDIF relies on identifying and segmenting large blood vessels, such as segments of the aorta or femoral arteries, to estimate the whole-blood time-activity curve for individual patients. These vessels are chosen primarily due to their large size, which facilitates effective correction of partial volume effects (PVE) [Zanotti_Fregonara_2012]. However, in brain PET studies, the small size of cerebral vessels often results in significant PVE, substantially decreasing the accuracy and precision of the derived IDIFs [Zanotti_Fregonara_2012]. To address this challenge, numerous segmentation methods have been proposed, utilizing either PET images alone or PET images co-registered with MRI [litton1997input, chen1998noninvasive, wahl1999regions, liptrot2004cluster, parker2005graph, su2005quantification, mourik2008partial, fung2009multimodal, lee2012extraction, fung2013cerebral, sari2017estimation, khalighi2018image]. Statistical decomposition approaches have further been introduced to mitigate PVE and noise. For example, chen1998noninvasive modeled each voxel signal in the carotid region-of-interest (ROI) as a mixture of vascular and spillover activity:
@@
Subsequent work removed the need for invasive sampling. su2005quantification applied ICA to the early dynamic PET frames and automated segmentation using Gaussian thresholding and dilation, while parker2005graph incorporated PCA and graph-based Mumford-Shah segmentation to refine carotid delineation . More recent efforts further improved artery segmentation through multimodal imaging and advanced algorithms [fung2009multimodal, khalighi2018image, su2013noninvasive, sari2017estimation, fung2013cerebral].
@@ 2.2 Direct estimation of IDIF
Alternatively, IDIF can be estimated directly from voxel values in dynamic PET, avoiding explicit segmentation of the carotid arteries [naganawa2005extraction, naganawa2005omission, naganawa2008robust, b2006extraction, wang2006model]. Such approaches commonly rely on statistical decomposition techniques, particularly independent component analysis (ICA), to separate voxel signals into plasma and tissue components. For example, Naganawa et al. proposed the following model to describe voxel values in dynamic PET images [naganawa2005extraction, naganawa2005omission]:
@@
where $s_{p}(q)$ and $s_{t}(q)$ represent the plasma and tissue contributions of the voxel $q$, respectively, while $c_{p}(t)$ and $c_{t}(t)$ denote the uncalibrated IDIF and the tissue time activity curve (tTAC) at time $t$, respectively. When expressed in matrix form, Eqn. 2 transforms into $\mathbf{X}=\mathbf{CS}$, where a modified ICA method known as EPICA decomposes the dynamic PET data matrix $\mathbf{X}$ into component matrices $\mathbf{C}$ and $\mathbf{S}$. The plasma-related component in $\mathbf{S}$ provides the uncalibrated IDIF. Similarly, Bodvarsson et al. employed non-negative matrix factorization (NMF) to derive the IDIF [b2006extraction]. Later, naganawa2008robust advanced the intersectional searching algorithm (ISA) [wang2006model], developing EPISA, which estimates a time-integrated IDIF directly from PET images for Logan graphical analysis [logan1990graphical], using tTACs obtained from intensity-based clustering. A key limitation of decomposition-based methods is that they often cannot inherently determine the scale or sign of estimated components, requiring calibration with a single blood sample, typically taken at the AIF peak.
@@ 3 Methods
[…]
@@
To achieve our goal, we introduce a novel DNN framework. This network is built on top of ViT [dosovitskiy2021an], enhanced with additional modules customized to process time-sequenced image volumes, such as dynamic PET scans. The output of ViT consists of a set of parameters that define the shapes of basis functions, which are then used to model the DLIFs.
@@ 3.1.1 Vision Transformer
Transformers, initially developed for natural language processing tasks [vaswani2017attention], have demonstrated significant potential in computer vision tasks [dosovitskiy2021an, liu2021swin, han2022survey]. Following these successes, they have been increasingly applied in the processing of medical images [li2023transforming]. A notable advantage of Transformers is their scalability [zhai2022scaling, liu2022convnet] and their capacity to capture long-range relationships between parts of the input. The decision to choose ViT over convolutional neural networks (ConvNets) for our application of AIF estimation is based on clear reasoning. ConvNets typically adopt small convolution kernels, usually ranging in size from 3 to 7, which inherently assume strong correlations between adjacent points (i.e., inductive bias). This presents two drawbacks for our application. First, AIF estimation often involves comparing spatially distant voxels, as seen in ISA [wang2006model] and EPISA [naganawa2008robust] described in the Introduction section. Localized convolution kernels in ConvNets are not as effective in capturing these distant relationships, which only become implicit as the network layers deepen. Second, for precise estimation across the temporal dimension, it is preferable to process the entire temporal data directly, rather than applying small convolution kernels across time points. It is important to acknowledge recent advancements that show that ConvNets can be optimized with much larger kernels [ding2022scaling, liu2023more] or advanced convolutional operations [liu2022convnet], potentially matching the capabilities of Transformers. However, optimizing the network architecture to balance Transformers and ConvNets is beyond the scope of our study. In this work, we have benchmarked the proposed network against commonly used DNNs, with further details and results presented in the Results section.
The proposed network is built on the foundation of ViT [dosovitskiy2021an], which was originally designed for 2D image processing. Recognizing the limitations of ViT in this regard, we have expanded its functionality to effectively handle 3D image volumes. The architecture of the neural network is depicted in Fig. 2, and the details regarding the ViT and the associated attention mechanism are described in A
@@ 3.1.2 Estimation head
In the original ViT [dosovitskiy2021an], the classification head is attached to the class token for the final prediction. In a similar fashion, our model attaches different estimation heads to the estimation token, $I^{\text{spatial}}\in\mathbb{R}^{1\times D}$, as shown in Fig. 2. Since this token engages with all others during self-attention, it is capable of learning the global information necessary for AIF estimation. The estimation heads are responsible for producing either a direct AIF estimation, which has an output dimension of $1\times T$, or a set of parameters to combine the basis functions. The latter is discussed in the following section.